6. Blood plasma has a pH range of 7.35 - 7.45. From the midpoint of this range calculate the following:

a) [H3O+]

b) [OH-]

c) pOH
1
Expert's answer
2020-10-29T05:25:23-0400

pH midpoint = 7.4

a)

pH = - lg[H3O+], so:

[H3O+] = 10-pH = 10-7.4 = 4 * 10-8 mol/L

b)

[H3O+] * [OH-] = 10-14, so:

[OH-] = 10-14 / [H3O+] = 2.5 * 10-7 mol/L

c)

pOH = 14 - pH = 14 - 7.4 = 6.6